find four geometric means between 1/8 and 4
Answer by stanbon(75887)

find four geometric means between 1/8 and 4
-----
a(1) = 1/8
a(5) = 4
--------
Equation:
a(5) = (1/8)r^4
4 = (1/8)r^4
r^4 = 32
r^4 = 2^5
r = 2*2^(1/4)
------
Ans: (1/8),(1/8)*2^5/4,(1/8)2^10/4,(1/8)2^15/4,(1/8)2^(20/4) - 4
=====================
Cheers,
Stan H.
--------------